Please help me solve this problem:

There is only one positive integer which is exactly twice the sum of its digits. Find this two digit number and show how you got your work.

Answer by 303795(602)

Trial and error sounds good (with a bit of extra thinking on the side)
If the number is twice the sum of its digits then the number must be even. Lets try 12.
12. 1 + 2 = 3 but 12 is not twice 3.
Using that idea the numbers have to total at least 7 so double 7 would be 14.
The two digit number with digits totalling 7 would be 16.
16. 1 + 6 = 7 but 16 is not twice 7 but we are getting closer (next even number to try is 18)
18. 1 + 8 = 9 and 18 is twice 9
So the number is 18.